# An Extension of the Adiabatic Theorem

### v1

@@ VI Overlaps in Post-Quench Spectrum
Figure 10: Quench within the floating phase. The phase diagram in panel 10 shows the starting and ending points of the quench. The quench from $i$ to $f$ is referred to as “forward”. Adapted from Karrasch et al. [14].
@@
Figure 11: Quench within the paramagnetic phase showing ambiguous behaviour.
The result appears ambiguous: the backward quench supports the conjecture’s statement, whereas the forward result does not. In the upper panel of Fig. 11, the largest peak is not located at the ground state energy but at higher energies in the spectrum. In fact, the heights of three peaks exceed that of the ground state overlap. A possible however not certain explanation for this observation could be finite-size shifted phase boundaries that are shown in Fig. 6.
To demonstrate that the above results are not a generic feature of any quenched system, but rather provide support for the conjecture, we show some quenches across phase transitions in Appendix IX. For such quenches, the largest peak is likely to occur at higher energy levels. Appendix Figures 12, 13, and 14 show quenches across various phase boundaries of the ANNNI model that were carried out along the depicted trajectories.
In contrast to the quenches within the same magnetic phase, the maximum overlap does not align with the ground state energy. Exceptions to this observation arise for quenches between the anti-phase and floating phase as can be seen in Appendix Fig. 14. The finite-size shifted phase diagrams in Fig. 6 suggest that the anti-phase is extended to significantly higher values of $h$ for the finite systems studied and thus provide a possible explanation for why the anti-phase-floating phase transition is not traceable.
@@ VII Conclusion
This work aimed to investigate the relationship between eigenstates of quenched Hamiltonians, specifically testing the validity of the conjecture formulated in Section II.
We conducted calculations in the TFIM and the ANNNI model. For the former, which is analytically solvable through Jordan-Wigner transformation and Bogoliubov rotation, we could establish a general analytical proof of the conjecture independent of specific quench points. For the ANNNI model, we could not attain a complete analytical proof as the model is not exactly solvable. Instead, we verified the conjecture in a special case where exact expressions for the pre- and post-quench ground states are available. To extend the analysis beyond this special case, we employed numerical methods.
The numerical analysis comprises ED and the Lanczos method to compute the Hamiltonian’s full spectrum or relevant parts of it. In the TFIM, numerical findings support the conjecture for all cases. Numerical results for the ANNNI model also mostly support the conjecture. While some violations could successfully be attributed to finite-size effects, for others it is not sure whether they can be explained with finite-size shifted phase boundaries. Employing other methods for quantifying finite-size effects may offer a more comprehensive explanation for the observed violations.
Overall, the findings cautiously affirm the conjecture, suggesting it as an extension of the adiabatic theorem by Born and Fock (1928)[10] to maximally non-adiabatic changes. We acknowledge that counterexamples can be constructed, at least for systems without a continuous spectrum. However, it remains unclear whether the conjecture holds universally for all systems with a continuous spectrum in the thermodynamic limit or if its validity must be restricted to a specific class of Hamiltonians or to quench end points that are relatively close to the quench start point. Further investigation is needed to determine the precise conditions under which the conjecture holds. Nevertheless, our results strongly suggest that a significant class of Hamiltonians satisfies the conjecture, making it worthwhile to explore the conditions under which it remains valid.
@@ VIII Acknowledgement
We thank Dirk Schuricht for his valuable input, which contributed significantly to this paper. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) through FOR 5522 (Project-ID 499180199).

### v2

@@ V Numerical Analysis of the ANNNI model
Figure 10: Quench within the floating phase. The phase diagram in (a) shows the starting and ending points of the quench. The quench from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
Within the phase where the quench starts, all points should ideally be displayed in the brightest color for the conjecture to hold true. Any darker color within that phase indicates a violation of the conjecture, as it signifies that the maximal overlap is not provided by the ground state overlap in such cases.
For all tested quench configurations within the ferromagnetic and floating phases and antiphase, the numerical data consistently support the conjecture (Fig. 2). However, for prequench ground states in the paramagnetic phase, the numerical results are partly inconclusive.
Figure 3 shows a starting point at $\Delta_{i}=-0.5$} successively approaching the quadruple point. When the quench starts far from this multiphase point, the numerical results support the conjecture. However, as the quench starting point moves closer to the quadruple point, violations of the conjecture begin to appear close to the Ising phase transition (Fig. 3). Figure 5 explores whether these violations could be caused by finite-size effects, as a shifted position of the phase boundaries for smaller systems is a known phenomenon in the ANNNI model [35].
The concept of a “phase transition” is not well-defined in finite systems, particularly when the considered systems are very small. Strictly speaking, the term applies to only systems in the thermodynamic limit, where phase transitions are characterized by a discontinuity in an order parameter or by nonanalytic behavior of derivatives of the free energy. As the Ising transition is a second-order phase transition, the free energy remains continuous in the thermodynamic limit, while higher derivatives become nonanalytic at the critical point. At finite sizes, however, the partition function is a finite sum of analytic functions, so the free energy and all its derivatives are analytic. Consequently, the singular features observed in the thermodynamic limit are rounded: Discontinuities become smooth crossovers, and divergences are replaced by finite-size peaks that sharpen and shift with increasing system size.
As mentioned in Sec. II, the fidelity susceptibility can be used as an indicator of quantum phase transitions. Thus, its finite-size version is a good candidate for locating the finite-size phase boundaries. For this purpose we make use of the fidelity susceptibility $\chi_{F}$ that occurs in the expansion of the finite-size fidelity (3).
By analyzing the parameter-dependent fidelity susceptibility we find that the location of the finite-size phase boundary generally approaches that of the infinite-size phase transition asymptotically. The technique is applied in two directions, given that the phase diagram is two-dimensional: The shift in the $h$ direction is determined while keeping $\Delta$ constant and vice versa.
The “remodeled” phase boundaries for a system of $L=16$ spins are depicted in Fig. 6. We are aware that the fidelity susceptibility is not a reliable measure for tracking phase transitions near the floating phase, as it typically exhibits peak revivals that depend on the system size (cf. [40]).
Some violations, such as those illustrated in Figs. 3(d) and 4(c), remain inexplicable by the adjusted positions of the phase transitions. However, other instances, such as the quenches depicted in Figs. 4(a), 5(a), and 5(b), can be accounted for by these revised phase transition locations.
VI Overlaps in the Post-Quench Spectrum
For quenches within the same phase, the conjecture states that the largest overlap with the initial ground state should be that with the final ground state. Consequently, when one plots the overlaps against the postquench Hamiltonian spectrum, the largest peak should align with the lowest energy state. In Figs. 7-14, we shift the postquench spectrum for this purpose such that the ground state energy is located at $0.0$ and all excitations energies are positive.
Figure 7(a) shows quenches within the paramagnetic phase, carried out along the depicted trajectories. The direction labeled “forward” denotes a quench from $i$ to $f$, while “backward” indicates a quench from $f$ to ii. Quenches within the floating phase, antiphase, and ferromagnetic phase are presented below using the same labeling scheme.
The results in Figs. 7-14 in the Appendix were obtained employing exact diagonalization. They support the conjecture, as the largest peak consistently occurs at the ground state energy.
This rendering of the results allows for a reinvestigation of the quench depicted in Fig. 4(c), which fails to support the conjecture as points near the Ising transition exhibit a larger overlap than the ground state overlap. A parameter combination for the quench starting at the same point as in Fig. 4(c) and ending in the darker region close to the phase boundary is used for the result illustrated in Fig. 11.
The result appears ambiguous: The backward quench supports the conjecture’s statement, whereas the forward result does not. In the top panel of Fig. 11(b), the largest peak is not located at the ground state energy but at higher energies in the spectrum. In fact, the heights of three peaks exceed that of the ground state overlap. A possible, although not certain explanation for this observation could be the finite-size shifted phase boundaries shown in Fig. 6.
To demonstrate that the above results are not a generic feature of any quenched system but, rather, provide support for the conjecture, we show some quenches across phase transitions in the Appendix. For such quenches, the largest peak is likely to occur at higher energy levels. Figures 12, 13, and 14 show quenches across various phase boundaries of the ANNNI model that were carried out along the depicted trajectories.
In contrast to the quenches within the same magnetic phase, the maximum overlap does not align with the ground state energy. Exceptions to this observation arise for quenches between the antiphase and floating phase as can be seen in Fig. 14 in the Appendix. The finite-size shifted phase diagrams in Fig. 6 suggest that the antiphase is extended to significantly higher values of $h$ for the finite systems studied and thus provide a possible explanation for why the antiphase–floating phase transition is not traceable.
@@ VI Overlaps in the Post-Quench Spectrum
Figure 11: Quench within the paramagnetic phase showing ambiguous behavior.
@@ VII Conclusion
This work aimed to investigate the relationship between eigenstates of quenched Hamiltonians, specifically testing the validity of the conjecture formulated in Sec. II.
We conducted calculations in the TFIM and the ANNNI model. For the former, which is analytically solvable through Jordan-Wigner transformation and Bogoliubov rotation, we were able to establish a general analytical proof of the conjecture independent of specific quench points. For the ANNNI model, we could not attain a complete analytical proof because the model is not exactly solvable. Instead, we verified the conjecture in a special case where exact expressions for the pre- and postquench ground states are available. To extend the analysis beyond this special case, we employed numerical methods.
The numerical analysis comprises ED and the Lanczos method to compute the Hamiltonian’s full spectrum or relevant parts of it. In the TFIM, numerical findings support the conjecture for all cases. Numerical results for the ANNNI model also mostly support the conjecture. While some violations could successfully be attributed to finite-size effects, for others whether they can be explained with finite-size shifted phase boundaries is not certain. Employing other methods for quantifying finite-size effects may offer a more comprehensive explanation for the observed violations.
Overall, the findings cautiously affirm the conjecture, suggesting it is an extension of the adiabatic theorem by Born and Fock[1] to maximally nonadiabatic changes. We acknowledge that counterexamples can be constructed, at least for systems without a continuous spectrum. However, whether the conjecture holds universally for all systems with a continuous spectrum in the thermodynamic limit or whether its validity must be restricted to a specific class of Hamiltonians or to quench end points that are relatively close to the quench start point remains unclear. Further investigation is needed to determine the precise conditions under which the conjecture holds. Nevertheless, our results strongly suggest that a significant class of Hamiltonians satisfies the conjecture, making it worthwhile to explore the conditions under which it remains valid.
@@ VIII Acknowledgement
We thank D. Schuricht for his valuable input, which contributed significantly to this paper. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Grant No. 499180199 (Project T1) via FOR 5522.
IX Data Availability
The data that support the findings of this article are openly available [41].
X Appendix
Figures 12, 13, and 14 show quenches across various phase boundaries of the ANNNI model that were carried out along the depicted trajectories.
[…]
Figure 12: Quenches across the paramagnetic–ferromagnetic phase transition. The phase diagram in (a) illustrates the quench trajectories. The quench direction from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
[…]
Figure 13: Quenches across the floating phase–paramagnetic phase transition. The phase diagram in (a) illustrates the quench trajectories. The quench direction from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
[…]
Figure 14: Quenches across the floating phase–antiphase transition. The phase diagram in (a) illustrates the quench trajectories. Adapted from Karrasch and Schuricht [32].